# An Attempt to Unraveling Token Prediction Refinement and Identifying Essential Layers of Large Language Models

### v1

@@ Abstract
This research aims to unravel how large language models (LLMs) iteratively refine token predictions (or, in a general sense, vector predictions). We utilized a logit lens technique to analyze the model’s token predictions derived from intermediate representations. Specifically, we focused on how LLMs access and use information from input contexts, and how positioning of relevant information affects the model’s token prediction refinement process. Our findings for multi-document question answering task, by varying input context lengths (the number of documents), using GPT-2, revealed that the number of layers between the first layer that the model predicted next tokens correctly and the later layers that the model finalized its correct predictions, as a function of the position of relevant information (i.e., placing the relevant one at the beginning, middle, or end of the input context), has a nearly inverted U shape. We found that the gap between these two layers, on average, diminishes when relevant information is positioned at the beginning or end of the input context, suggesting that the model requires more refinements when processing longer contexts with relevant information situated in the middle, and highlighting which layers are essential for determining the correct output. Our analysis provides insights about how token predictions are distributed across different conditions, and establishes important connections to existing hypotheses and previous findings in AI safety research and development.

### v2

# Unraveling Token Prediction Refinement and Identifying Essential Layers in Language Models

@@ Abstract
This research aims to unravel how large language models (LLMs) iteratively refine token predictions through internal processing. We utilized a logit lens technique to analyze the model’s token predictions derived from intermediate representations. Specifically, we focused on (1) how LLMs access and utilize information from input contexts, and (2) how positioning of relevant information affects the model’s token prediction refinement process. On a multi-document question answering task with varying input context lengths, we found that the depth of prediction refinement (defined as the number of intermediate layers an LLM uses to transition from an initial correct token prediction to its final, stable correct output), as a function of the position of relevant information, exhibits an approximately inverted U-shaped curve. We also found that the gap between these two layers, on average, diminishes when relevant information is positioned at the beginning or end of the input context. This suggested that the model requires more refinements when processing longer contexts with relevant information situated in the middle. Furthermore, our findings indicate that not all layers are equally essential for determining final correct outputs. Our analysis provides insights into how token predictions are distributed across different conditions, and establishes important connections to existing hypotheses and previous findings in AI safety research and development.